A coin with a diameter of 2.1 cm is dropped on edge onto a horizontal surface. The coin starts out with an initial angular speed of 15 rad/s and rolls in a straight line without slipping. The rotation slows with an angular acceleration of magnitude 1.7 rad/s$^2.
a) What is the initial linear velocity of the coin
b)What is the magnitude of the linear acceleration?
c)How far does the coin roll before it stops?
1
Expert's answer
2020-09-24T11:08:27-0400

solution

given data-

initial angular velocity("\omega_0" )=15 rad/s

radius of coin(R)=2.1/2 =1.05 cm

angular acceleration("\alpha" )=1.7 rad/s^2


(a)initial linear velocity can be given as for pure rotation


"v=\omega _0R"

then

"v=15\times1.05\times10^{-2}=" "0.16 m\/s"


(b) linear acceleration can be written as

"a=\alpha R"


then

"a=1.7\times1.05\times10^{-2}=0.02m/s^2"


(c)

by applying equation of motion for rotation


"\omega^2=\omega_0^2-2\alpha \theta"


when coin stop then "\omega=0"

then

"\theta=\frac{\omega_0^2}{2\alpha}=\frac{15^2}{2\times 1.7}=66.18rad"


distance traveled by coin


"s=\theta R \\s=66.18\times 1.05\times 10^{-2}\\s=0.6949m\\ and \\s=0.7m"


so distance rolled by coin before stop is 0.7 m.